OpenStudy (anonymous):

Suppose an isosceles triangle ABC has A = PI/4 and b = c = 4. What is the length of a^2? a) 4^(2)(2-SQR2)) b) 4^(2)(SQR2) c) 4^(2)(2+SQR2) d) 4^(2)(SQR2 -2)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

a² = b² + c² -2bc cos A = 4² + 4² - 2(4)(4)* cos 45 = 16+16-32 (1/√2) = 16+16-16(√2) = 4² {2- √2 }

OpenStudy (anonymous):

Thanks :)